OpenStudy (anonymous):

the zeroes of the poly.p(x)=2x^-9 are:- (a) 9/2,-9/2 (b) -3/2,3/2 (c) 2/9,-2/9 (d) root 2/3,-root 2/3

OpenStudy (apoorvk):

well, equate the polynomial to zero, and solve for 'x'.

OpenStudy (anonymous):

will u plzz do ittt plzzzz

OpenStudy (apoorvk):

is that a 2x^2? I am assuming so.. correct me if am wrong. so,\[2x^2-9=0\] solve this, can't you?

OpenStudy (apoorvk):

\[2x^2 = 9\] \[=> x^2 = \frac9 2\] now what?

OpenStudy (anonymous):

wht is dis???????